# Discovering universal temperature regulation dynamics in animals

### v1

@@ 3.5 Discovered model is consistent with Hourglass and Threshold Hypothesis
Directly relating the hidden state $y(t)$ to a particular metabolite or physiological state is impossible without direct measurements. However, we can computationally explore possible mechanisms and hypothesize about the minimal mechanism of regulation. In the simplest case, we can imagine the hidden state as a molecule that is depleting over time as envisioned by Ruf, Giroud, and Geiser [8] (see Figure 3C top). However, by introducing a symmetry transformation into the hidden state of the form $\tilde{y}=-y+\psi$, we can generate dynamics that become negative (see Figure 3C bottom) or correspond to an accumulating, potentially toxic molecule [8]. The first parameterization of the scaling symmetry may correspond to a scenario where the threshold acts on the difference between multiple molecules rather than an individual entity ($\psi=0.1$). The second parameterization of the scaling symmetry reflects a fact mentioned in [8]: there is no way to distinguish between an accumulating metabolite and a depleting metabolite as the driver of interbout arousals ($\psi=0.75$). This class of symmetry transformations does not affect the fit to the measured state, so we cannot distinguish which of these model variants and corresponding metabolic scenarios are correct. Whether representing a single molecule or interactions between multiple molecules, the effective state acts as an Hourglass and Threshold-type mechanism.
3.6 Thermoregulatory transitions as a change of stability
@@
The Arctic ground squirrel thermophysiology model we discovered (Eqs. 3-4) can also describe seasonal transitions in the thermophysiology. In the summer, the Arctic ground squirrel is homeothermic, meaning its body temperature remains near 37 C with minor fluctuations around that physiological fixed point induced by activity and sleep. During hibernation, the Arctic ground squirrel’s body temperature exhibits oscillatory spiking dynamics, known biologically as poikilothermy. Physiologically, the organism cannot remain solely at 37 C or the low state (near or below 0 degrees Celsius) during hibernation, but rather must oscillate between the two states to survive, so if any physiological fixed points corresponding to high-temperature or low-temperature states exist, they must be unstable (or non-attractive). Mathematically, we expect the stability of the model’s high-temperature fixed point to vary in a seasonal manner, and the model’s low-temperature fixed point should always be unstable, as staying at a low temperature without oscillations at any time of year would be fatal for the organism. To confirm this, we perform a stability analysis of our model.
@@
To perform a stability analysis, we can nondimensionalize Eqs. (3)-(4)) with non-essential terms $\theta_{A}$=$\theta_{B}$=$\theta_{C}$=$\theta_{D}$=0 and take $\gamma\approx\xi$, which is true of the parameters estimated during validation (see Section 6.3). The simplified nondimensionalized model has the form
{dimensionless body $\{\text{dimensionless body temperature}\}~{}~{}\dot{\hat{x}}$ $=\hat{x}\left(\hat{x}-\gamma\hat{y}\right)\left(1-\hat{x}\right)+\iota,\hat{x}(0)=\hat{x}_{0},$ (7)
@@
{dimensionless unknown physiological $\{\text{dimensionless unknown physiological state}\}~{}~{}\dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (8)
@@
This model has three fixed points: $(\hat{x},\hat{y})=(0,0)$, $(1,-1)$, and $(1,1)$ when $\iota=0.$  [^4] It is reasonable to take $\iota=0$ to examine fixed points sufficiently far away from the origin, as the higher order terms will locally dominate $\iota$, as $\iota<<1$, but this is not reasonable near the origin where $\iota$ dominates. So under the $\iota=0$ approximation, we cannot comment on the stability of the fixed point at the origin as our assumption does not hold. Additionally, note that the fixed point $(\hat{x},\hat{y})=(1,-1)$ is unstable for all $\gamma$ and $\kappa$. A linear stability analysis of Eqs. (7)-(8) when $\iota=0$ shows that $(\hat{x},\hat{y})=(1,1)$, which corresponds to the high-temperature state, is stable for $10<\gamma<$ 1 and is an unstable saddle for $\gamma>1$. For $\gamma=1$, linear stability analysis is indeterminate. Biologically, the thermoregulatory transitions between homeothermy and poikilothermy can be understood as seasonal variation in $\gamma$, perhaps driven by the circannual clock, which corresponds mathematically to a seasonal change of stability of the fixed point $(\hat{x},\hat{y})=(1,1)$, the high-temperature fixed point. While the Arctic ground squirrel is homeothermic during the summer ($\gamma<1$), the thermophysiological system stays near $(\hat{x},\hat{y})=(1,1)$ with minor fluctuations. During hibernation ($\gamma>1$), $(\hat{x},\hat{y})=(1,1)$ is unstable, and the system snaps to the limit cycle, which describes the poikilothermic regime. Thus, the stability of the point near (1,1) is controlled by $\gamma$ in a seasonal manner as expected.
@@
Notice that the second inequality will be dominated by the second term, $-8\gamma\kappa$, as $\iota<<1$ in our simulations. This leads to complex eigenvalues with positive real parts and, therefore, an unstable equilibrium near the origin. As expected, the stability of $(\hat{x},\hat{y})=(1,1)$ varies in a seasonal manner and $(\hat{x},\hat{y})=\left(\sqrt{\dfrac{-\iota}{(1-\gamma)}},\sqrt{\dfrac{-\iota}{(1-\gamma)}}\right)$ is always unstable in the parameter regime described by the Inequalities (9)-(10).
3.7 Thermophysiological organization across species
Many species make thermoregulatory switches similar to the Arctic ground squirrel (Figure 4A), but often the timescale or amplitude of the oscillations are different (Figure 4B). It is currently unknown if the same thermophysiological system is conserved across species or if different species have fundamentally different systems. Here, we show how our discovered core model can be modified to capture body temperature patterns from a wide range of species, suggesting the possibility of a unified mechanism across species that can respond to environmental signals.
Here, we show the possibility of a conserved thermophysiological mechanism by modifying our discovered core model (Equations (7)-(8)) to respond to environmental information. To determine how to add environmental information to our models, we first examine the dimensionality of the thermophysiological system across species using time delay embedding. We find that the dimensionality of the underlying thermophysiological systems for the elephant shrew, noisy miner, and black bear are relatively low dimensional across species (all systems appear to be three dimensional, see Figure 4C and Sections 5.2-5.3). To approximate the extra dimensionality, we add environmental information to our core model. Mathematically, these take the form of trigonometric forcing functions. We find that our modified models can qualitatively describe the body temperature patterns of “daily-heterotherm” species (the noisy miner), an ”intermediate” species (the elephant shrew), and a large ”hibernator” (the black bear) (Figure 4B and D).
Figure 4: We modify our discovered physiological model of Arctic ground squirrel to include environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, a shrew, and a bear, which also dynamically modulate body temperature. This suggests that a universal, environmentally sensitive mechanism could regulate body temperature across a diverse range of species. A: Photographs of an Arctic ground squirrel, an elephant shrew, a noisy miner, and a black bear. All images taken from Wikipedia. B: Temperature recordings from an Arctic ground squirrel [9], an elephant shrew [16], a noisy miner [31], and a black bear [14]. C: Phase space reconstructions of underlying thermophysiological systems. These systems are 2-to-3 dimensional. D: Model simulations of body temperature across species (blue) qualitatively match the observed body temperature data in B. Inferred dynamics of a hidden physiological state (pink). E: Environmental information interacting with the core model for each species.
4 Discussion
Here, we have used a blend of automated model discovery and classical dynamical systems theory to differentiate between two physiological mechanisms hypothesized to drive interbout arousals during small-mammalian hibernation and comment on the possible organization of thermophysiological systems across species using heterothermic endothermy. We find evidence that interbout arousals are driven by an Hourglass and Threshold-type mechanism, as hypothesized in [8]. We modified our discovered, core model to include environmental signals and found that we could qualitatively match body temperature recordings from a variety of species, including a bird, a shrew, and a black bear (Figure 4). This leads to a new hypothesis: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to environmental signals.
Beyond the algorithmic design, we made several practical modeling choices. We included monomials up to order three in our library during model selection. We drew inspiration from classical dynamical systems models from neuroscience, which typically include terms up to third degree, such as the FitzHugh-Nagumo Model [23, 24]. Models with higher-order terms, such as quartic and quintic terms, may work as well or better but would be more computationally challenging to select. Another choice was which term to add to the selected models to constrain the period of the limit cycle. We chose to add a small constant to the body temperature equation as the simplest option, but other small linear terms would also regulate the period length of the limit cycle. In future work, we wish to understand if our models can be transformed into Líenard systems or if we have discovered any candidates from a broader and distinct class of oscillatory models.
With regards to the biological interpretation, our effective model of Arctic ground squirrel thermophysiology provides information that can be helpful during the molecular search for the biological entity that causes interbout arousals. In particular, our model supports the hypothesis that the biological entity varies with the same period as the interbout arousals. However, our model cannot differentiate between an accumulating or depleting entity, or if the driver is a single entity or an amalgamation of multiple entities. In the qualitative modeling, we did not examine all possible ways to add environmental information. We demonstrate the existence of one version of a core model and a suite of extended, environmentally driven models, but we do not expect uniqueness. Similarly, it is important to note that the hidden state in our models of thermophysiology across species may not correspond to the same underlying biological entity in each species. We simply claim that the thermophysiological systems may have core components that interact in the same fashion across species, whether due to genetic conservation or convergence.
Despite these limitations and caveats, our analysis provides mechanistic insight into hibernation thermophysiology and puts forward a hypothesis in thermal biology: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to environmental signals. Our analysis starts to describe how physiological systems are affected and controlled by the external environment, and as more physiology is uncovered, we can further refine our mechanistic models. In the future, we hope to further understand which features of the external environment, such as photoperiod, ambient temperature (see Figure 2 in [33]), magnetic fields, social interactions, and food availability influence thermophysiology and physiology more broadly in potentially nonlinear and seasonally-varying ways.

### v2

@@ 3.4 Discovered model is consistent with Hourglass and Threshold Hypothesis
Directly relating the hidden state $y(t)$ to a particular metabolite or physiological state is impossible without direct measurements. However, we can computationally explore possible mechanisms and hypothesize about the minimal mechanism of regulation. In the simplest case, we can imagine the hidden state as a molecule that is depleting over time as envisioned by Ruf, Giroud, and Geiser [8] (see Figure 3C top). However, by introducing a symmetry transformation into the hidden state of the form $\tilde{y}=-y+\psi$, we can generate dynamics that become negative (see Figure 3C bottom) or correspond to an accumulating, potentially toxic molecule [8]. If we take $\psi=0.1$, then the scaling symmetry may correspond to a physiological scenario where the threshold acts on the difference between multiple molecules rather than an individual entity (See Figure 3 C bottom). If we take $\psi=0.75$, then $\tilde{y}(t)$ stays completely positive and inverts the dynamics of $y(t)$ (top, Figure 3C). This parameterization of the scaling symmetry reflects a fact mentioned in [8]: there is no way to distinguish between an accumulating metabolite and a depleting metabolite as the driver of interbout arousals. This class of symmetry transformations does not affect the fit to the measured state, so we cannot distinguish which of these model variants and corresponding metabolic scenarios are correct. Whether representing a single molecule or interactions between multiple molecules, the effective state acts as an Hourglass and Threshold-type mechanism.
3.5 Towards a multi-scale model of seasonal physiology
Small hibernators exhibit clear seasonal changes in physiology (see Figure 1). In the summer, the Arctic ground squirrel is homeothermic, with minor fluxuations induced by activity and sleep around a physiological fixed point near 37 C, but in hibernation, the organism cannot remain solely at 37 C or the low state (near or below 0 degrees Celsius), but must oscillate between the two states to survive. Despite being found using sparse model-selection on only the data during hibernation, our discovered model captures these dynamics, has parameters that have clear physiological meaning, and connects with the known biology of the circannual clock [35, 36].
Consider the nondimensionalization of Eqs. (3)-(4) with with non-essential terms $\theta_{A}$$\theta_{B}$$\theta_{C}$$\theta_{D}$=0,
$\{\text{dimensionless body temperature}\}~~\dot{\tilde{x}}$ $=\omega\tilde{x}^{2}-\tilde{x}^{3}-\nu(\tilde{x}\tilde{y}-\tilde{x}^{2}\tilde{y})+\zeta,$ $\tilde{x}(0)=\tilde{x}_{0},$ (5)
$\{\text{dimensionless unknown physiological state}\}~~\dot{\tilde{y}}$ $=\rho(\tilde{x}^{2}-\tilde{y}^{2}),$ $\tilde{y}(0)=\tilde{y}_{0}.$ (6)
Each of the parameters has a clear physiological interpretation. The parameter $\omega$ controls the amplitude of the oscillations during hibernation. From laboratory experiments, it is known that hibernators attenuate the amplitude of the oscillations in response to environmental factors [37]. The parameter $\nu$ controls the seasonally varying coupling between body temperature and the unknown physiological state. $\nu$ is the parameter that controls the transitions from homeothermy to poikilothermy and vice versa. Biologically, $\nu$ can be interpreted as a summary output of the circannual clock (see below). The parameters $\zeta$ and $\rho$ influence the period of oscillations during hibernation. $\rho$ controls the scale separation between body temperature and the unknown physiological state and influences the period of oscillations. Biologically, we can interpret $\zeta$ and $\rho$ as modulating the speed of the hourglass and the location of the threshold, perhaps in an environmentally sensitive manner [37].
The system undergoes a supercritical Hopf bifurcation as a function of $\nu$, and we can find where this bifurcation occurs in $\omega$-$\nu$-$\zeta$-$\rho$ space.  [^4] It is straightforward to analytically find where this bifurcation occurs in $\omega$-$\nu$-$\zeta$-$\rho$ space. This curve divides parameter space into qualitatively different behaviors and has the form $\nu^{4}-8\nu^{3}\zeta+2\nu^{3}\rho-4\nu^{3}\omega-4\nu^{2}\rho^{2}+4\nu^{2}\omega^{2}-8\nu\rho^{3}+16\nu\rho^{2}\omega-8\nu\rho\omega^{2}-2\nu^{3}+$ (7) $36\nu^{2}\zeta-6\nu^{2}\rho+9\nu^{2}\omega+12\nu\rho\omega-12\nu\omega^{2}+8\rho^{3}-12\rho^{2}\omega+4\omega^{3}-54\nu\zeta+27\zeta=0.$ The curve is visualized in $\omega$-$\nu$ space for fixed values of $\zeta$ and $\rho$ (see Figure 4 A). The transition from a regulated body temperature to an oscillating body temperature during hibernation is associated with a change of stability in the system (see Section 5) and traversing a boundary in parameter space (see Figure 4). Similarly, the exit from hibernation and reversion to a regulated body temperature exhibited in the summer is associated with the opposite change of stability and movement in parameter space (see Section 5 and Figure 4A). Assuming $\nu$ is time-dependent, we can find an example trajectory of $\nu(t)$ that allows the system to capture the broad features of seasonal hibernation physiology (Figure 4 B and C).
Figure 4: The beginnings of a multi-scale, neuro-physiological understanding of seasonal physiology. A: We find the location of the supercritical Hopf bifurcation in $\omega$-$\nu$-$\zeta$-$\rho$ space and plot Eq. 7 for $\zeta=0.001$ and $\rho=0.1$. This curve divides parameter space into two qualitatively different model behaviors associated with distinct thermoregulatory modes exhibited by the Arctic ground squirrel: a stable fixed point, which corresponds to a regulated body temperature exhibited by the squirrel in the summer, and oscillatory dynamics exhibited during hibernation. Seasonal thermoregulatory changes are associated with the system crossing the boundary that separates fixed point behavior and oscillatory dynamics. B: Biologically, $\nu(t)$ may correspond to a summary output of the circannual clock, a neural mechanism to track seasonal change and generate seasonal decisions. An example trajectory of $\nu(t)$ that allows Eqs. (5)-(6) to qualitatively capture the body temperature dynamics observed throughout the annual cycle (see Figure 1, Left). C: The full seasonal physiological dynamics of Eqs. (5)-(17) as $\nu(t)$ varies as in B. Dimensionless body temperature, $\tilde{x}$, and the dimensionless unknown physiological state, $\tilde{y}$, are shown in blue and purple, respectively.
The dynamics of $\nu(t)$ shown in Figure 4 B is qualitatively consistent with our current understanding of the circannual clock. The circannual clock is a neural mechanism that tracks the passage of the year and generates mammalian seasonal decisions, such as hibernation timing, migration timing, mating decisions, and seasonal coat change in organisms like the snowshoe hare and the Arctic fox. The long time scale of these behaviors has made a molecular description of the circannual clock elusive, but some is known about the mechanistic underpinnings of the clock. The clock is thought to be located in the pars tuberalis in the Anterior pituitary gland, and seasonal decisions are thought to be controlled by a population of pars-tuberalis-specific thyrothropic cells, which exhibit a bistable switch in gene expression [36]. This change causes downstream endocrinological changes that form the basis of seasonal decision-making. At the tissue level, the theorized dynamics of the thyrothropic cells qualitatively resemble the trajectory of $\nu$ over time [36]. We interpret $\nu$ as a summary output of the circannual clock, and this model as the beginnings of a multi-scale model of seasonal physiology that connects the neurological and physiological scales.
3.6 Thermophysiological organization across species
Many species make thermoregulatory switches similar to the Arctic ground squirrel (Figure 5A), but often the timescale or amplitude of the oscillations is different (Figure 5B). It is currently unknown if the same thermophysiological system is conserved across species or if different species have fundamentally different systems. Here, we show how our discovered core model can be modified to capture body temperature patterns from a wide range of species, suggesting the possibility of a unified mechanism across species that can respond to environmental signals.
Here, we show the possibility of a conserved thermophysiological mechanism by modifying our discovered core model (Equations (10)-(11)) to respond to environmental information. To determine how to add environmental information to our models, we first examine the dimensionality of the thermophysiological system across species using time delay embedding. We find that the dimensionality of the underlying thermophysiological systems for the elephant shrew, noisy miner, and black bear is relatively low-dimensional across species (all systems appear to be three-dimensional, see Figure 5C and Sections 6.2-6.3). To approximate the extra dimensionality, we add environmental information to our core model. Mathematically, these take the form of trigonometric forcing functions. We find that our modified models can qualitatively describe the body temperature patterns of “daily-heterotherm” species (the noisy miner), an ”intermediate” species (the elephant shrew), and a large ”hibernator” (the black bear) (Figure 5B and D).
Figure 5: We modify our discovered physiological model of Arctic ground squirrel to include environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, an elephant shrew, and a bear, which also dynamically modulate body temperature. This suggests that a universal, environmentally sensitive mechanism could regulate body temperature across a diverse range of species. A: Photographs of an Arctic ground squirrel, an elephant shrew, a noisy miner, and a black bear. All images taken from Wikipedia. B: Temperature recordings from an Arctic ground squirrel [9], an elephant shrew [23], a noisy miner [39], and a black bear [21]. C: Phase space reconstructions of underlying thermophysiological systems. These systems are 2-to 3-dimensional. D: Model simulations of body temperature across species (blue) qualitatively match the observed body temperature data in B. Inferred dynamics of a hidden physiological state (pink). E: Environmental information interacting with the core model for each species.
4 Discussion
Here, we have used a blend of automated model discovery and classical dynamical systems theory to differentiate between two physiological mechanisms hypothesized to drive interbout arousals during small-mammalian hibernation and comment on the possible organization of thermophysiological systems across species using heterothermic endothermy. We find evidence that interbout arousals are driven by an Hourglass and Threshold-type mechanism, as hypothesized in [8] and work towards a multi-scale understanding of seasonal physiological decision making through a bifurcation analysis (see Section 3.5). We modified our discovered core model to include environmental signals and found that we could qualitatively match body temperature recordings from a variety of species, including a bird, an elephant shrew, and a black bear (Figure 5). This leads to a new hypothesis: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to environmental signals.
Beyond the algorithmic design, we made several practical modeling choices. We included monomials up to order three in our library during model selection. We drew inspiration from classical dynamical systems models from neuroscience, which typically include terms up to third degree, such as the FitzHugh-Nagumo Model [14, 15]. Models with higher-order terms, such as quartic and quintic terms, may work as well or better but would be more computationally challenging to select. Another choice was which term to add to the selected models to constrain the period of the limit cycle. We chose to add a small constant to the body temperature equation as the simplest option, but other small linear terms would also regulate the period length of the limit cycle. In future work, we wish to understand if our models can be transformed into Líenard systems or if we have discovered any candidates from a broader and distinct class of oscillatory models. In the qualitative modeling, we did not examine all possible ways to add environmental information. We demonstrate the existence of one version of a core model and a suite of extended, environmentally driven models, but we do not expect uniqueness. Similarly, it is important to note that the hidden state in our models of thermophysiology across species may not correspond to the same underlying biological entity in each species. We simply claim that the thermophysiological systems may have core components that interact in the same fashion across species, whether due to genetic conservation or convergence.
Despite these limitations and caveats, our discovered macro-scale physiological model can be used to guide the ongoing search for molecular mechanisms that underpin small mammalian hibernation physiology based on the dynamics of the hidden state. However, our model cannot differentiate between an accumulating or depleting entity, or whether the driver is a single entity or an amalgamation of multiple entities. In practice, the causative agent may take the form of a single metabolite, the ratio of key metabolites, a difference between key metabolites, or something else entirely. Our model of hibernation physiology would be validated if the currently unknown causative agent oscillates with the same period as body temperature dynamics during hibernation and is constant during summer, when the organism exhibits a stable body temperature, or invalidated if not. A key point is that the causative agent may not be low in the summer. It may be relatively high, but we expect it to be constant. It is important to note that our qualitative modeling presented in see Section 3.5 captures only the broad features of the seasonal body temperature time series and was not intended to model the “flickering” dynamics that occur at the entrance and exit of hibernation [12]. In future work, we will modify the model to better capture these dynamics, but this was not the intention of this modeling effort.
We discover a new data-driven physiological model in the style of the Fitzhugh-Nagumo model [14, 15], the NPZ model [16], and the Brusselator model [17, 18, 19], which distill the incredible complexity and retain only the essential features of the underlying system. We used our discovered model to provide new insights into hibernation physiology, posit the beginnings of a multi-scale model of hibernation that spans the neurological and physiological scales, and put forward a hypothesis in thermal biology: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to environmental signals. Our analysis starts to describe how physiological systems are affected and controlled by the external environment, and as more physiology is uncovered, we can further refine our models. In the future, we hope to further understand which features of the external environment, such as photoperiod, ambient temperature (see Figure 2 in [41]), magnetic fields, social interactions, and food availability, influence thermophysiology and physiology more broadly in potentially nonlinear and seasonally varying ways.
5 Appendix
@@
The Arctic ground squirrel thermophysiology model we discovered (Eqs. 3-4) can also describe seasonal transitions between summer homeothermy and winter poikilothermy. In the summer, the Arctic ground squirrel is homeothermic, with minor fluctuations induced by activity and sleep around a physiological fixed point near 37 C. During hibernation, the organism cannot remain solely at 37 C or the low state (near or below 0 $\circ$ C), but must oscillate between the two states to survive. Therefore, any physiological fixed points corresponding to high-temperature or low-temperature states must be unstable (or non-attractive). Mathematically, we expect the stability of the model’s high-temperature fixed point to transition from stable to unstable in the fall and back to stable in the spring. The model’s low-temperature fixed point should always be unstable, as staying at a low temperature without oscillations at any time of year would be fatal for the organism. To confirm this, we perform a stability analysis of our model.
@@
To perform a stability analysis, we perform a second nondimensionalization on Eqs. (3)-(4)) with non-essential terms $\theta_{A}$$\theta_{B}$$\theta_{C}$$\theta_{D}$=0. The simplified nondimensionalized model has the form
$\{\text{dimensionless body temperature}\}~~\dot{\hat{x}}$ $=\hat{x}^{2}-\hat{x}^{3}-\gamma\hat{x}\hat{y}+\xi\hat{x}^{2}\hat{y}+\iota,\hat{x}(0)=\hat{x}_{0},$ (8)
@@
$\{\text{dimensionless unknown physiological state}\}~~\dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (9)
If we assume $\gamma\approx\xi$, which is true of the parameters estimated during validation (see Section 7.3), Eqs. (8)-(9) has the following simplified form
$\{\text{dimensionless body temperature}\}~~\dot{\hat{x}}$ $=\hat{x}\left(\hat{x}-\gamma\hat{y}\right)\left(1-\hat{x}\right)+\iota,\hat{x}(0)=\hat{x}_{0},$ (10)
$\{\text{dimensionless unknown physiological state}\}~~\dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (11)
@@
The model parameters of Eqs. (10)-(11) has interpretable physiological meaning. $\gamma$ describes the seasonal coupling between body temperature and the unknown physiological state. The parameters $\iota$ and $\kappa$ influence the period of oscillations during hibernation. Eqs. (10)-(11) has three fixed points: $(\hat{x},\hat{y})=(0,0)$, (1,1)(1,-1), and $(1,1)$ when $\iota=0.$ [^5] It is reasonable to take $\iota=0$ to examine fixed points sufficiently far away from the origin, as the higher order terms will locally dominate $\iota$, as $\iota<<1$, but this is not reasonable near the origin where $\iota$ dominates. So under the $\iota=0$ approximation, we cannot comment on the stability of the fixed point at the origin, as our assumption does not hold. Additionally, note that the fixed point $(\hat{x},\hat{y})=(1,-1)$ is unstable for all $\gamma$ and $\kappa$. A linear stability analysis of Eqs. (10)-(11) when $\iota=0$ shows that $(\hat{x},\hat{y})=(1,1)$, which corresponds to the high-temperature state, is stable for $0<\gamma<1$ and is an unstable saddle for $\gamma>1$. For $\gamma=1$, linear stability analysis is indeterminate. Biologically, the thermoregulatory transitions between homeothermy and poikilothermy can be understood as seasonal variation in $\gamma$, perhaps driven by the circannual clock, which corresponds mathematically to a seasonal change of stability of the fixed point $(\hat{x},\hat{y})=(1,1)$, the high-temperature fixed point. While the Arctic ground squirrel is homeothermic during the summer ($\gamma<1$), the thermophysiological system stays near $(\hat{x},\hat{y})=(1,1)$ with minor fluctuations. During hibernation ($\gamma>1$), $(\hat{x},\hat{y})=(1,1)$ is unstable, and the system snaps to the limit cycle, which describes the poikilothermic regime. Thus, the stability of the point near (1,1) is controlled by $\gamma$ in a seasonal manner as expected. This seasonal transition and change of stability are shown in Figure 4 for the same model with a different choice of nondimensionalization.
@@
Notice that the second inequality will be dominated by the second term, $8\gamma\kappa$, as $\iota<<1$ in our simulations. This leads to complex eigenvalues with positive real parts and, therefore, an unstable equilibrium near the origin. As expected, the stability of $(\hat{x},\hat{y})=(1,1)$ varies in a seasonal manner and $(\hat{x},\hat{y})=\left(\sqrt{\dfrac{-\iota}{(1-\gamma)}},\sqrt{\dfrac{-\iota}{(1-\gamma)}}\right)$ is always unstable in the parameter regime described by the Inequalities (12)-(13).